### v1

# The Economics of Information Pollution in the Age of AI: A General Equilibrium Approach to Welfare, Measurement, and Policy

Abstract
The advent of Large Language Models (LLMs) represents a fundamental shock to the economics of information production. By asymmetrically collapsing the marginal cost of generating low-quality, synthetic content while leaving high-quality production costly, AI systematically incentivizes information pollution. This paper develops a general equilibrium framework to analyze this challenge. We model the strategic interactions among a monopolistic platform, profit-maximizing producers, and utility-maximizing consumers in a three-stage game. The core of our model is a production technology with differential elasticities of substitution ($\sigma_{L}>1>\sigma_{H}$), which formalizes the insight that AI is a substitute for labor in low-quality production but a complement in high-quality creation. We prove the existence of a unique ”Polluted Information Equilibrium” and demonstrate its inefficiency, which is driven by a threefold market failure: a production externality, a platform governance failure, and an information commons externality. Methodologically, we derive a theoretically-grounded Information Pollution Index (IPI) with endogenous welfare weights to measure ecosystem health. From a policy perspective, we show that a first-best outcome requires a portfolio of instruments targeting each failure. Finally, considering the challenges of deep uncertainty, we advocate for an adaptive governance framework where policy instruments are dynamically adjusted based on real-time IPI readings, offering a robust blueprint for regulating information markets in the age of AI.
Introduction
The rise of large language models (LLMs) marks a significant turning point in information economics, revolutionizing content production and profoundly impacting market efficiency, social welfare, and institutional stability. Unlike traditional information asymmetry, the challenge posed by LLMs lies in the near-zero marginal cost of producing low-quality content that appears plausible but lacks factual support or originality. This undermines traditional quality signals, as the production cost of high-quality information (which requires expertise and rigorous verification) remains high. This asymmetry shifts the relative production costs of high- and low-quality content in information markets, fostering a proliferation of low-quality content. Industry reports indicate that API costs for generative AI have fallen by over $90\%$ in two years, while the annual economic losses from disinformation exceed \$100\\$100 billion, highlighting the profound impact of this shift.
Given this, the paper addresses four central research questions. First, how does the asymmetric nature of AI technology reshape the equilibrium in the market for information? Second, what are the precise mechanisms of market failure in this new equilibrium, and how can their welfare consequences be characterized? Third, can we construct a theoretically-grounded, operational metric to measure the health of the information ecosystem in real time? Fourth, what is the optimal design of a multi-instrument policy framework that can correct these novel market failures in an environment of deep technological uncertainty?
To answer these questions, we develop a general-equilibrium model of the information market. Our central theoretical contribution lies in formalizing the asymmetric technological shock through production functions with differential elasticities of substitution, capturing that AI capital is a gross substitute for human labor in low-quality content production ($\sigma_{L}>1$) but a gross complement in high-quality creation ($\sigma_{H}<1$).
The core contribution of this study lies in theoretically demonstrating the uniqueness and inefficiency of the ”polluted information equilibrium” and explicitly identifying its triple market failure. Methodologically, we propose the ”Information Pollution Index” (IPI), a theoretically-grounded, multi-dimensional composite index derived directly from our model to quantify ecosystem health. Policy-wise, we derive an optimal multi-instrument framework, showing that achieving the first-best outcome requires a combination of tools like Pigouvian taxes, content provenance standards, and platform fiduciary obligations to simultaneously address these market failures. Recognizing implementation challenges under uncertainty, we further propose an adaptive governance framework that leverages the IPI as a real-time feedback signal.
The remainder of the paper is organized as follows. Section II lays out the theoretical framework. Section III formally defines and characterizes the Polluted Information Equilibrium, including its existence and inefficiency. Section IV develops the Information Pollution Index (IPI) as a theory-based metric. Section V solves for the optimal policy portfolio. Section VI discusses robustness and the challenges of deep uncertainty, motivating an adaptive governance approach. Section VII concludes with policy implications and directions for future research.
Literature Review
The economic theoretical basis and impact mechanism of information pollution
The digital age, characterized by unprecedented information flow and technological advancements, has given rise to a novel and complex market failure: information pollution. While its theoretical foundations can be traced back to Akerlof’s (Akerlof 1970) ”lemon market” model, its manifestation in the contemporary digital environment presents unique, multi-dimensional complexities. On one hand, false information spreads rapidly through the topological structure of social networks (Vosoughi, Roy, and Aral 2018), exhibiting more significant network externalities (Allcott and Gentzkow 2017) and being continuously reinforced by cognitive biases (such as the truth illusion effect revealed by Pennycook &\& Rand (Pennycook and Rand 2019)). On the other hand, information pollution causes multidimensional economic distortions, such as (Allcott et al. 2020), (Durante, Pinotti, and Tesei 2019) and (Bursztyn et al. 2020).
If we consider it from the perspective of pollution, environmental economics has two important tools for dealing with pollution: Pigouvian taxes and tradable permits. A Pigovian tax is an economic tool designed to correct market failures caused by negative externalities. In recent years, some scholars have begun to explore the application of Pigovian taxes to digital platform governance. Dranove et al. (Dranove et al. 2014) proposed the idea of levying an ”attention tax” on online advertising to alleviate the problem of information overload. Cabral et al. (Cabral et al. 2021) further developed this idea and proposed a differentiated tax mechanism for low-quality content. Zuboff (Zuboff 2019) proposed the concept of ”information extraction tax” in The Age of Surveillance Capitalism, its theoretical framework is also inspiring for information pollution governance. Tradable permits, first proposed by Dales (Dales 1968) and based on the property rights theory of Coase (Coase 1960), that imposes a tradable hard cap (cap) on the total amount of pollutants that can be emitted by a particular industry or economy. Montgomery (Montgomery 1972) demonstrated that, under certain conditions, tradable permits can achieve pollution control objectives with minimal costs. Although there are some obstacles to shifting this perspective to information pollution, some scholars have put forward some relevant proposals. Parker &\& Van Alstyne (Parker and Van Alstyne 2021) proposed the transaction of information quality points between platforms to solve the problem of information asymmetry in the platform market. In a report written for the European Commission, Crémer et al. (Crémer, de Montjoye, and Schweitzer 2019) proposed the concept of ”content quotas”, similar to a tradable license mechanism.
The impact of AI on the information market: governance challenges and multidisciplinary integration research
There are significant challenges in applying traditional economic governance tools to information pollution governance. In the information field, these approaches are more likely to raise questions about fairness, effectiveness, and potential conflicts with the right to freedom of expression, and may lead to censorship or discriminatory law enforcement. In practical terms, it is also difficult for a tradable license system to clearly define the ”upper limit” of information pollution. Without a strong definition, it will be very difficult to implement such a system fairly and effectively, which may distort competition and consolidate the market dominance of large platforms. Existing understandings also lack the need for long-term analysis. All of this highlights the need for comprehensive impact assessments and flexible regulatory frameworks. This paper contributes to and bridges three major strands of literature. First, we extend the classic literature on information economics. We recast the “market for lemons” ((Akerlof 1970)) for the digital age, and we explore how AI undermines traditional signaling mechanisms ((Spence 1973)). Second, we contribute to the literature on platform economics and industrial organization ((Rochet and Tirole 2004); (Hagiu and Wright 2015)). Our model is the first to formally integrate a sophisticated production-side analysis of AI’s impact into a platform intermediation framework. Third, we build upon the burgeoning literature on the economics of AI ((Agrawal, Gans, and Goldfarb 2018); (Acemoglu and Restrepo 2019)). Unlike most research on the labor market, our focus is on its impact on the quality of output in one of society’s most critical markets: the information market.
Theoretical Framework
This section develops the core theoretical framework of the paper. We first specify the economic environment, agents, and types of information goods. Second, we characterize the asymmetric production technology for content, which serves as the fundamental driver of pollution. Finally, we formalize the strategic interactions among the platform, producers, and consumers as a three-stage Stackelberg game, outlining its equilibrium characteristics and the sources of market failure.
The Economic Environment
Agents.
Our economic system consists of four types of rational agents, each with well-defined objectives and constraints:
  • Content Producers: A continuum of profit-maximizing producers indexed by $i\in[0,1]$. Each producer $i$ is endowed with a heterogeneous productivity parameter $A_{j,i}$ and faces factor prices $(r,w)$, where $r$ is the rental rate of AI capital and $w$ is the wage rate for high-skilled human labor.
  • Monopolistic Platform: Acting as the intermediary, the platform wields market power and serves a dual role. As a market organizer, it sets the content moderation intensity $m\in[0,1]$. As an attention allocator, it determines the exposure weights for different content types via an algorithmic amplification vector $\bm{\gamma}=(\gamma_{H},\gamma_{L})$.
  • Consumers: A representative group of consumers with heterogeneous verification costs $k_{i}$, drawn from a distribution $F(k)$ with support $0,\bar{k}]$. Consumers observe a noisy signal $s\in\{H,L\}$ of content quality and decide whether to pay $k_{i}$ to perfectly ascertain it.
  • Social Planner: A theoretical benchmark character whose objective is to maximize social welfare. The planner allows us to define the Pareto efficient allocation and quantify the welfare loss from market failures.
Information Goods.
Two fundamentally different types of information goods exist in the market:
  • High-Quality Content ($Q_{H}$): Information products characterized by factual accuracy, cognitive value that improves consumer decision-making, and production complexity requiring expertise and rigorous verification.
  • Low-Quality Content ($Q_{L}$): Information products that may be misleading or false, generate negative externalities by harming consumer welfare and social trust, and are simple to produce via templated or automated means.
This classification is based not on subjective judgment but on the objective impact of content on consumer decisions and social welfare.
Production Technology: The Dynamics of Pollution
CES Production Function Specification.
The core mechanism of our model lies in the production technology. For content of type $j\in\{H,L\}$, producers utilize two inputs: AI capital ($K_{AI}$) and high-skilled human labor ($L_{H}$). The production function is of the Constant Elasticity of Substitution (CES) form:
$Q_{j}=A_{j}\bigl{[}\delta_{j}K_{AI}^{\rho_{j}}+(1-\delta_{j})L_{H}^{\rho_{j}}\bigr{]}^{1/\rho_{j}}$ (1)
where $A_{j}$ is the total factor productivity, $\delta_{j}$ is the distribution parameter, and $\rho_{j}$ determines the elasticity of substitution, $\sigma_{j}=1/(1-\rho_{j})$.
Core Assumption: Technological Asymmetry.
Our theory is built upon the following key assumption:
Assumption 1 (Technological Asymmetry).
$\sigma_{L}>1>\sigma_{H}>0$ (2)
This assumption has profound economic implications:
  • For low-quality content ($\sigma_{L}>1$), AI capital and human labor are gross substitutes. This reflects the standardized nature of producing such content, where AI can efficiently generate text that is syntactically correct but lacks deep verification.
  • For high-quality content ($\sigma_{H}<1$), AI capital and human labor are gross complements. This captures the complexity of creating such content, where AI tools can augment research and writing but cannot replace human creativity, critical analysis, and ethical judgment.
Cost Asymmetry.
A direct corollary of this assumption is the asymmetric effect of technological progress on production costs.
Proposition 1 (Cost Asymmetry).
A decrease in the cost of AI capital, $r$, has a larger cost-reducing effect on low-quality content than on high-quality content. Formally:
$\left|\frac{\partial\log c_{L}}{\partial\log r}\right|>\left|\frac{\partial\log c_{H}}{\partial\log r}\right|$ (3)
Intuition: As AI becomes cheaper, the production cost of ”lemons” ($Q_{L}$) falls more sharply than that of ”peaches” ($Q_{H}$). This asymmetric cost advantage is the fundamental supply-side driver of information pollution. (For a formal proof, see Appendix A.1).
The Platform-Producer-Consumer Game
We model the market interaction as a three-stage sequential game, following the logic of Stackelberg competition.
Stage I: Platform Strategy Selection.
The platform, as the Stackelberg leader, first chooses its governance policy mix $(m,\gamma_{H},\gamma_{L})$ to maximize its profit function $\Pi_{P}$:
$\Pi_{P}=\theta\rho\bigl{[}\gamma_{H}Q_{H}^{S}+\gamma_{L}(1-m)Q_{L}^{S}\bigr{]}-C_{m}(m)$ (4)
where $\theta$ is the platform’s revenue share, $\rho$ is the average ad revenue per amplified unit of content, $Q_{j}^{S}$ is the anticipated producer supply, and $C_{m}(m)$ is a convex moderation cost function.
Stage II: Producer Supply Decision.
Observing the platform’s policy, producers choose their optimal supply. The profit per unit of content $j$ is:
$\pi_{j}=(1-\theta)\rho\gamma_{j}-c_{j}(r,w)$ (5)
Heterogeneity in producer productivity $A_{j,i}$ allows for the derivation of a smooth industry supply curve $Q_{j}^{S}(\gamma_{H},\gamma_{L})$, which satisfies $\partial Q_{j}^{S}/\partial\gamma_{j}>0$ (direct incentive effect) and $\partial Q_{j}^{S}/\partial\gamma_{k}<0$ for $k\neq j$ (resource competition effect).
Stage III: Consumer Verification Decision.
Consumers observe a noisy signal $s\in\{H,L\}$ and decide whether to pay their verification cost $k_{i}$. The signal precision, $\pi(s=H|q=H)=\pi(\rho^{\prime},V^{*})$, is endogenous to the health of the information ecosystem, where $\rho^{\prime}$ is the effective pollution density and $V^{*}$ is the aggregate verification rate. This function satisfies $\partial\pi/\partial\rho^{\prime}<0$ (pollution dilutes signal quality) and $\partial\pi/\partial V^{*}>0$ (verification creates positive externalities). The aggregate verification rate $V^{*}$ must satisfy a fixed-point condition, capturing the collective action nature of the verification problem.
Equilibrium and The Threefold Market Failure
Definition 1 (Polluted Information Equilibrium).
A Subgame Perfect Nash Equilibrium (SPNE) is a strategy profile and outcome combination $\{(m^{*},\bm{\gamma}^{*}),(Q_{H}^{*},Q_{L}^{*}),V^{*},\pi^{*}\}$ such that the decisions of the platform, producers, and consumers are mutually optimal, and beliefs are consistent with outcomes.
Theorem 1 (Existence of Equilibrium).
Under standard regularity conditions, a Polluted Information Equilibrium exists.
(For a proof sketch, see Appendix A.2).
Theorem 2 (Market Failure).
The Polluted Information Equilibrium is Pareto inefficient, driven by three interacting market failures:
  1. 1.
    Production Externality: Producers of low-quality content do not internalize the negative social effects of their output. The social marginal cost exceeds the private marginal cost:
    $MC_{\text{social}}=MC_{\text{private}}+\frac{\partial}{\partial Q_{L}}\bigl{[}d(Q_{L}^{\prime})+\lambda\frac{\partial T}{\partial Q_{L}}\bigr{]}$ (6)
    where $d(\cdot)$ is the direct consumer harm function and $T$ is the stock of social trust with shadow price $\lambda$.
  2. 2.
    Platform Governance Failure: The platform’s objective function is misaligned with social welfare. An engagement bias assumption implies that the platform may find it profitable to set $m^{*}<m_{SO}$ and $\gamma_{L}^{*}>\gamma_{L,SO}$.
  3. 3.
    Information Commons Externality: Verification is a public good, but consumers are not compensated for its full social benefit, leading to under-investment. The social marginal benefit exceeds the private marginal benefit:
    $MB_{\text{social}}=MB_{\text{private}}+\frac{\partial\pi}{\partial V}\cdot(\text{Welfare gain for others})$ (7)
Proposition 2 (Paradox of AI Progress).
A decrease in the cost of AI capital, $r$, leads to:
  1. 1.
    an increase in the supply of low-quality content ($\partial Q_{L}^{*}/\partial r<0$),
  2. 2.
    a rise in effective pollution density ($\partial\rho^{\prime*}/\partial r<0$),
  3. 3.
    a decline in decentralized social welfare ($\partial W^{*}/\partial r>0$).
(For a proof sketch, see Appendix A.3).
The Information Pollution Index
Based on the market equilibrium characteristics revealed by the preceding theoretical model, we construct an Information Pollution Index (IPI) that is endogenous to the social welfare function. This section aims to build a solid bridge from abstract theory to observable measurement, ensuring the index possesses both a rigorous theoretical foundation and practical applicability for policy.
Theoretical Foundation and Axiomatic Properties of the IPI
Theoretical Construction.
Definition 2 (Information Pollution Index).
The Information Pollution Index at time $t$, denoted $IPI(t)$, is defined as a linear combination of its four theoretical dimensions:
$IPI(t)=\sum_{j=1}^{4}w_{j}^{*}(t)\cdot I_{j}(t)$ (8)
where the weight $w_{j}^{*}(t)$ is endogenously determined by the marginal impact of each dimension on social welfare. This weight reflects the marginal rate of social aversion to each type of harm at a given equilibrium state $E^{*}(t)$:
$w_{j}^{*}(t)=\frac{\left|\frac{\partial W}{\partial I_{j}}\right|_{E^{*}(t)}}{\sum_{k=1}^{4}\left|\frac{\partial W}{\partial I_{k}}\right|_{E^{*}(t)}}$ (9)
Economic Properties.
An effective economic index should satisfy several desirable axiomatic properties. We show that the IPI meets the following key criteria:
Property 1 (Welfare Monotonicity).
The IPI is strictly negatively correlated with social welfare ($\partial W/\partial IPI<0$). This property establishes the index as a ”social welfare thermometer,” ensuring that its rise unambiguously indicates a deterioration in Pareto efficiency.
Property 2 (Decomposability).
Changes in the index can be clearly attributed to its constituent dimensions, as $\partial IPI/\partial I_{j}=w_{j}^{*}$. This allows policymakers not only to observe the overall level of pollution but also to diagnose the core drivers of the problem for targeted interventions.
Property 3 (Policy Sensitivity).
The total effect of any policy intervention on the IPI can be decomposed into a direct effect on the level of each dimension and an indirect effect on the weights via changes in the equilibrium structure (a Lucas-critique-style effect). This property provides a rigorous framework for counterfactual policy evaluation.
Precise Definitions and Economic Interpretations of the Four Dimensions
The IPI is composed of four dimensions, each capturing a distinct facet of the harm caused by information pollution.
First, we define Effective Pollution Density ($I_{1}$), which measures not a simple ratio of content volume, but the effective share of pollution in the consumer attention market. It is given by:
$I_{1}(t)=\frac{\gamma_{L}^{*}(t)(1-m^{*}(t))Q_{L}^{*}(t)}{\gamma_{H}^{*}(t)Q_{H}^{*}(t)+\gamma_{L}^{*}(t)(1-m^{*}(t))Q_{L}^{*}(t)}$ (10)
This dimension captures a core economic reality: low-quality content, amplified by platform algorithms ($\gamma_{L}^{*}$) in pursuit of engagement and escaping moderation ($1-m^{*}$), has an impact on finite mental bandwidth that far exceeds its absolute quantity. It is a direct characterization of pollution’s prevalence and salience.
Second, we define Social Welfare Deadweight Loss ($I_{2}$), which monetizes the harm of information pollution by measuring the economic cost borne by society due to the threefold market failure. Conceptually, it is equivalent to the Harberger’s Triangle in public finance, quantifying the net efficiency loss from resource misallocation. Its formal definition is:
$I_{2}(t)=\frac{W^{SO}-W^{*}(t)}{W^{SO}-W^{\text{min}}}$ (11)
This dimension answers the key question: ”How much value does our society lose due to information pollution?”
Third, we consider the Decay of the Trust Commons ($I_{3}$), which treats ”trust” as a depletable stock of social capital. Information pollution acts as a corrosive agent, eroding the foundation of market transactions and social cooperation. The indicator captures the long-term, cumulative, and path-dependent harm of pollution, defined as:
$I_{3}(t)=\frac{T^{\max}-T(t)}{T^{\max}}$ (12)
where the stock of trust $T(t)$ follows the dynamic path $\frac{dT(t)}{dt}=-\delta I_{1}(t)\text{Flow}(t)+\xi R(t)-\mu T(t)$. Steady-state analysis suggests that persistent high pollution can push society into a low-trust poverty trap.
Fourth, we define Asymmetric Technology Risk ($I_{4}$), which measures the core technological driver of pollution—a technological arms race between content generation and detection. It is a forward-looking risk indicator of the ecosystem’s systemic vulnerability, defined as:
$I_{4}(t)=\frac{1}{2}\left[1+\tanh\left(\frac{\log\left(\frac{\text{Cap}_{\text{gen}}(t)}{\text{Cap}_{\text{det}}(t)}\right)-\mu_{\text{tech}}}{\sigma_{\text{tech}}}\right)\right]$ (13)
When the capability of generation technology (”offense”) far outstrips that of detection technology (”defense”), the marginal cost of producing ”lemons” falls, systemically exacerbating the other three dimensions of harm. Graphical framework reference picture 1.
Figure 1: Information Pollution Index (IPI) Four-Dimensional Radar Chart Across Policy Scenarios. This radar chart displays the four dimensions of the Information Pollution Index across different scenarios: (I) Effective Pollution Density, $(I_{2})$ Social Welfare Deadweight Loss, $(I_{3})$ Decay of Trust Commons, and $(I_{4})$ Asymmetric Technology Risk. Each axis represents normalized values from 0 (best) to 1 (worst). The baseline scenario reflects current market conditions without intervention. Al cost reduction demonstrates the pollution amplification effect of technological progress. Policy intervention shows the effectiveness of the proposed multi-instrument framework. The first-best optimum represents the theoretical benchmark for comparison. Larger areas indicate higher pollution levels and welfare losses.
From Theory to Measurement: The Core Proxy Indicator System
Since the theoretical dimensions are not directly observable, we design a core proxy indicator for each, aiming to unify theoretical validity with empirical tractability.
Our proxy for Effective Pollution Density ($I_{1}$) is the Weighted Exposure Pollution Rate. This is calculated as the ratio of impressions from low-quality content to total impressions:
$\hat{I}_{1}=\frac{\sum_{i}\mathbf{1}_{\{i\text{ is low-quality}\}}\cdot\text{Impressions}_{i}}{\sum_{i}\text{Impressions}_{i}}$ (14)
We choose ”impressions” over ”clicks” as the weight because it more closely reflects the initial supply of pollution, being less affected by the endogenous variable of user discernment. This indicator directly measures the ”market share” of low-quality content in the platform’s attention allocation mechanism, serving as a direct counterpart to the theoretical dimension $I_{1}$.
To proxy for Social Welfare Deadweight Loss ($I_{2}$), we propose the Weighted Harm Feedback Rate. A simple complaint rate cannot distinguish the welfare loss from ”clickbait” versus ”financial fraud.” By weighting different types of negative user feedback by their potential harm severity, this indicator measures the intensity of harm as revealed by consumer revealed preferences. It is thus a more effective ”pain signal” for welfare loss, calculated as:
$\hat{I}_{2}=\frac{\sum_{j}\text{HarmType}_{j}\cdot\text{Weight}_{j}\cdot\text{FeedbackCount}_{j}}{\text{Total Impressions}}$ (15)
For the Decay of the Trust Commons ($I_{3}$), we propose the Trust-Related Churn Gap via Causal Inference. The decay of trust ultimately manifests as users ”voting with their feet.” To address attribution challenges, we employ quasi-experimental methods (e.g., DiD, RDD) to compare user cohorts randomly or quasi-randomly exposed to different levels of pollution. The resulting difference in their churn rates, normalized by a baseline, quantifies the net effect of trust erosion on economic outcomes, consistent with the gold standard of modern empirical economics. The calculation is:
$\hat{I}_{3}=\frac{\text{Churn}_{\text{high-exposure}}-\text{Churn}_{\text{low-exposure}}}{\text{Churn}_{\text{baseline}}}$ (16)
Finally, to proxy for Asymmetric Technology Risk ($I_{4}$), we use the Public Benchmark Detection Accuracy Gap. This indicator quantifies the ”offense-defense gap” via standardized adversarial benchmarking, measuring the performance of the best ”shield” (current detection systems) against the sharpest ”spear” (latest generative models). Changes in this indicator correspond directly to changes in the key exogenous technology parameters of our theoretical model, bridging the theory of endogenous technological change with observable reality. It is calculated as:
$\hat{I}_{4}=1-\frac{\text{Accuracy}_{\text{on new models}}}{\text{Accuracy}_{\text{on baseline data}}}$ (17)
Composite Index Construction and Application
After obtaining standardized time-series data for the four core proxy indicators, we advocate for a hybrid approach—combining theoretical deduction, empirical estimation, and expert judgment—to determine their weights. A robust data quality assurance framework is also essential. The resulting IPI can be used for causal impact evaluation in academic research and as a practical tool for long-term ecosystem health monitoring by platforms and regulators, thus closing the loop between theoretical analysis, empirical monitoring, and policy intervention.
The Static Optimal Policy Portfolio: A Theoretical Benchmark
In an ideal world with complete information and a stable technological structure, the policymaker’s task is to design a set of instruments that target each of the three market failures respectively.
To correct the production externality, classic Pigouvian instruments can be employed. The optimal tax, $\tau_{L}^{*}$, levied directly on the output of low-quality content $Q_{L}$, should equal its marginal social damage:
$\tau_{L}^{*}=d^{\prime}(Q_{L,SO}^{\prime})(1-m_{SO})+\lambda^{*}\frac{\partial T}{\partial Q_{L}}$ (18)
where the first term represents the direct harm to consumers and the second term captures the damage to the stock of social trust (with shadow price $\lambda^{*}$).
To correct the information commons externality arising from under-investment in verification, policy can directly enhance the average precision of public signals, $\pi$, through mandatory content provenance standards. This increases the perceived disutility of low-quality content, thereby suppressing its production and dissemination throughout the game’s equilibrium path.
To correct the platform governance failure, policy can impose information fiduciary duties. This is formally equivalent to modifying the platform’s objective function to a weighted sum of its own profit and social welfare:
$\max_{m,\bm{\gamma}}\;(1-\alpha)\Pi_{P}+\alpha\bigl{[}v(Q_{H}^{\prime})-d(Q_{L}^{\prime})\bigr{]}$ (19)
where the intensity of the fiduciary duty, $\alpha\in[0,1]$, becomes a key regulatory choice variable.
Theorem 3 (First-Best Policy Portfolio).
In a static environment, a policy portfolio $(\tau_{L}^{*},\pi_{SO},\alpha^{*})$ consisting of an optimal Pigouvian tax, mandatory provenance standards, and information fiduciary duties can implement the first-best social optimum.
Table 1 summarizes this idealized mapping.
Table 1: The Triad of Market Failures and Corresponding Policy Instruments.
Market Failure Locus Consequence Instrument
Production Externality Producers Over-production of $Q_{L}$ $\tau_{L}$ / Permits
Info. Commons Ext. Consumers Under-verification Provenance
Platform Gov. Fail. Platform $\uparrow\gamma_{L},\downarrow m$ Fiduciary Duty
Dynamic Challenge: The Lucas Critique and Policy-Induced Innovation
However, the effectiveness of any static policy portfolio faces the fundamental challenge of the Lucas (1976) critique. Rational agents will strategically react to policy interventions, thereby altering the very economic structure upon which the policy was based. In our model, a sustained tax $\tau_{L}$ on low-quality content will lower the marginal return to innovations that enhance its production efficiency, $A_{L}$. Rational R&D efforts will thus be reallocated.
Proposition 3 (Policy-Induced Innovation Bias).
The imposition of $\tau_{L}>0$ induces a technology shift biased towards high-quality content, i.e., $\partial(A_{L}^{*}/A_{H}^{*})/\partial\tau_{L}<0$, where $A_{j}^{*}$ are endogenous productivity levels.
This dynamic feedback implies that any fixed ”optimal” tax rate $\tau_{L}^{*}$ will become suboptimal over time. The static policy blueprint is not only inefficient but also unreliable in a dynamic context.
Deep Uncertainty Challenge: Knightian Uncertainty and Robust Decision-Making
Beyond predictable endogenous dynamics, the long-term trajectory of AI is fraught with deep, unquantifiable uncertainty, i.e., Knightian uncertainty (Knight 1921). We lack a firm basis for assigning probabilities to future technological paradigms or ”black swan” events. In such a context, a rational social planner should adopt a more robust decision criterion than simple expected utility maximization, such as max-min expected utility (Gilboa and Schmeidler 1989):
$\max_{\mathcal{P}_{\text{policy}}}\;\min_{p\in\mathcal{P}}\;\mathbb{E}_{p}\bigl{[}W(\mathcal{P}_{\text{policy}})\bigr{]}$ (20)
The planner optimizes against the worst-case scenario within a set of plausible priors $\mathcal{P}$.
Proposition 4 (Optimality of the Precautionary Principle).
Under Knightian uncertainty, the optimal policy is inherently more precautionary, favoring higher taxes, stricter caps, and a greater reliance on ”fail-safe” instruments like content provenance standards.
The presence of Knightian uncertainty demands that policy design be resilient not only to known dynamics but also to unknown shocks.
A Synthesis: IPI-Centered Adaptive and Robust Governance
Facing the structural fragility revealed by the Lucas critique and Knightian uncertainty, an effective governance framework must transcend the static-optimality paradigm. We argue that the Information Pollution Index (IPI) constructed in the previous section is the core operational tool to bridge this gap from theory to practice.
The IPI, as a real-time, welfare-linked ”dashboard,” enables a feedback-based adaptive regulation. Policy instruments are no longer fixed constants but are governed by a state-contingent rule that adjusts based on the observed health of the ecosystem:
$\tau_{t}=\tau_{t-1}+\eta\left(\frac{IPI_{t-1}-IPI_{\text{target}}}{IPI_{\text{target}}}\right)$ (21)
Under this rule, the tax rate automatically tightens when the IPI exceeds its target and loosens otherwise. This mechanism endogenously responds to systemic drifts driven by technological innovation or market structure changes, thus partially addressing the Lucas critique. Furthermore, by continuously monitoring the IPI and its sub-dimensions, policymakers can achieve earlier detection of unforeseen risks (early signals of ”black swans”), enabling timely activation of contingency plans, which is the essence of robust decision-making.
In summary, confronting the governance challenges of the AI era requires abandoning the quest for a rigid, one-off policy blueprint. Instead, we must construct a multi-layered governance system. At its core is principles-based regulation (e.g., safety, transparency, accountability) to ensure long-term stability. Its method involves regulatory sandboxes for exploratory learning to navigate technological uncertainty. Its navigation system relies on real-time monitoring tools like the IPI to shift from reactive responses to anticipatory governance. Together, these three components form a resilient and adaptive ”governance vessel” capable of navigating the deep uncertainties of the digital ocean.
Experimental Validation
To validate our theoretical framework and the efficacy of the Information Pollution Index (IPI), we designed a comprehensive suite of experiments using an Agent-Based Model (ABM). By constructing a simulation environment that captures the three-party game among content producers, consumers, and a platform, we test our core theoretical predictions and assess the reliability of the IPI as a metric for information ecosystem health.
Experimental Design and Methodology
Simulation Model Architecture.
We construct an information economy simulation system with three classes of heterogeneous agents. Producer Agents, endowed with heterogeneous productivity parameters $(A_{H,i},A_{L,i})$, make content creation decisions based on expected profit maximization and a CES production function. Consumer Agents, characterized by heterogeneous verification costs $k_{i}$ and risk preferences, make consumption and verification decisions based on Bayesian updating and utility maximization, with externalities mediated through a social network. The Platform Agent, acting as a market intermediary, influences content distribution through algorithmic weights $(\gamma_{H},\gamma_{L})$ and moderation intensity $m$, aiming to maximize ad revenue under a user retention constraint.
Key Parameter Settings.
Based on our theoretical model, we set the following core parameters for the baseline simulation: elasticity of substitution for high-quality content $\sigma_{H}=0.75$ (complementarity) and for low-quality content $\sigma_{L}=1.5$ (substitutability); 100 producer agents and 300 consumer agents; baseline AI capital cost $r=1.0$ and labor cost $w=8.0$; and a platform revenue share of $\theta=0.25$.
Experiment 1: Comprehensive IPI Validation
This experiment was designed to thoroughly validate the theoretical effectiveness, measurement robustness, and predictive power of the IPI. The design included six sub-experiments: a baseline evolution test, a shock response test, a weight sensitivity analysis, a noise robustness test, a historical event detection simulation, and a cross-platform comparison.
The core findings strongly support the IPI’s validity. First, the IPI exhibits a strong negative correlation with social welfare (correlation coefficient of -0.839), confirming its theoretical validity as a welfare metric. Second, the index is highly sensitive to external shocks, showing an average increase of 37.5% during four simulated shocks, proving its ability to reflect real-time changes in ecosystem health. Third, the IPI is robust to different weighting schemes, with the IPI-welfare correlation remaining high (between 0.486 and 0.710) across six different configurations. Fourth, it demonstrates excellent resilience to measurement noise, with a measurement error of only 0.045 even at a 20% noise level. Finally, in a simulated fake news event, the IPI increased by 21.6%, showcasing its potential as an early warning tool. Our leading indicator analysis further reveals that the IPI has an optimal prediction window of 5 time steps, providing a valuable reaction window for policymakers.
Figure 2: Comprehensive experimental results for the validation of the Information Pollution Index (IPI). This figure summarizes the findings from Experiment 1, including baseline evolution, component breakdown, IPI-welfare correlation, shock response, and various robustness and sensitivity analyses.
Experiment 2: Theoretical Validation and Policy Analysis
Validation of the AI Progress Paradox.
We conducted a parameter sweep to systematically test the impact of AI cost ($r$) and the elasticity of substitution for low-quality content ($\sigma_{L}$) on the market equilibrium. The parameter space covered $r\in\{0.6,0.8,1.0,1.2,1.4\}$ and $\sigma_{L}\in\{1.2,1.4,1.6,1.8\}$. The key finding is a strong validation of the ”paradox of AI progress”: AI cost is positively correlated with social welfare (coefficient of 0.167) and strongly negatively correlated with pollution density (coefficient of -0.770). For instance, at an AI cost of 0.6, pollution density reached a high of 0.954, whereas it decreased to 0.568 when the cost was 1.4. This result provides robust computational support for our proposition that technological progress, due to its asymmetric impact, can paradoxically exacerbate information pollution and harm social welfare.
Analysis of Policy Interventions.
We designed six policy configurations to test the effectiveness of different intervention tools. Table 2 details the performance of key macroeconomic indicators under each policy.
Table 2: Policy Comparison Experiment Results.
Policy Scenario Welfare Pollution IPI Trust
1. Baseline 78.05 0.774 0.694 0.312
2. Pigouvian Tax ($\uparrow\theta$) 78.68 0.663 0.654 0.323
3. Subsidy ($\downarrow k_{max}$) 79.29 0.612 0.634 0.236
4. Joint Policy 79.82 0.753 0.636 0.272
5. Tech Intervention 79.34 0.596 0.622 0.277
6. Efficiency Boost 78.97 0.657 0.651 0.321
As shown in Table 2, all interventions improved upon the baseline across various dimensions. Notably, the Joint Policy achieved the highest social welfare (79.82), a 2.3% improvement over the baseline. The Tech Intervention was most effective at reducing pollution density (to 0.596) and the IPI (to 0.622). A key finding is that the effect of single instruments is limited; for example, the Pigouvian tax proxy, while beneficial, was outperformed by comprehensive strategies. This again validates our theoretical claim that a multi-tool policy portfolio is necessary to effectively address the threefold market failure.
Figure 3: Simulation results for the balanced Agent-Based Model, corresponding to Experiment 2. This figure illustrates the long-term evolution of key system-level indicators, the relationships between core theoretical variables (e.g., AI cost vs. pollution), the comparative effectiveness of different policy interventions, and analyses of system stability.
Discussion
Our simulation experiments provide substantial validation for the paper’s core theoretical contributions. They confirm the significance of asymmetric production technology, demonstrate the interplay of the threefold market failure mechanism, and show the convergence of the system to a suboptimal equilibrium with high pollution, consistent with our theoretical predictions. The experiments also establish the practical value of the IPI as a monitoring tool with early-warning capabilities, cross-context applicability, and utility for policy evaluation.
The results yield important policy implications: single-instrument policies are insufficient; a forward-looking regulatory framework is necessary to address the rapid evolution of AI; and policies must be adaptive, adjusting dynamically based on real-time data. While our simulation provides valuable insights, we acknowledge its limitations, such as model simplification and reliance on calibrated rather than empirically estimated parameters. Future research could enhance this framework by calibrating the model with real-world platform data, expanding the complexity of agent behaviors, and conducting longer-term dynamic analyses.
Appendix A Appendix A: AI pollution theory framework.
Figure 4: AI pollution theory framework.
Appendix B Appendix B: Model and Experimental Details
This appendix provides supplementary details on the Agent-Based Model (ABM) specification, parameter settings, experimental procedures, and results to ensure the reproducibility of our study.
B.1 Agent-Based Model Specification
The simulation is built upon the interaction of three agent types whose behavioral rules are designed to operationalize the core theoretical mechanisms.
Producer Agent Logic.
The production decision of producer $i$ between high-quality ($H$) and low-quality ($L$) content is governed by a softmax (logit) choice model based on expected profits:
$\text{Prob}(i\text{ chooses }H)=\frac{\exp(\beta\cdot\mathbb{E}[\pi_{H,i}])}{\exp(\beta\cdot\mathbb{E}[\pi_{H,i}])+\exp(\beta\cdot\mathbb{E}[\pi_{L,i}])}$ (22)
where $\mathbb{E}[\pi_{j,i}]$ is the expected profit for content type $j$ given agent $i$’s productivity, and $\beta$ is a rationality parameter that introduces bounded rationality.
Consumer Agent Logic.
A consumer $i$ with verification cost $k_{i}$ decides to verify a piece of content if the expected utility gain from resolving uncertainty exceeds the cost. The verification threshold $k_{i}^{*}$ is determined by:
$k_{i}^{*}=p_{i}(H|s)\cdot\Delta U_{H}+(1-p_{i}(H|s))\cdot\Delta U_{L}$ (23)
where $p_{i}(H|s)$ is the consumer’s posterior belief that the content is high-quality given signal $s$, and $\Delta U_{j}$ is the utility difference between informed and uninformed consumption of content type $j$.
Platform Agent Logic.
The platform agent adaptively adjusts its algorithmic amplification vector $\bm{\gamma}$ and moderation intensity $m$ based on market feedback. The rules follow a gradient-ascent logic, balancing revenue maximization with a user trust (retention) constraint:
$\gamma_{L,t+1}$ $=\gamma_{L,t}+\eta\left(\frac{\partial\Pi_{P}}{\partial\gamma_{L}}-\lambda\frac{\partial\text{Trust}}{\partial\gamma_{L}}\right)$ (24)
$m_{t+1}$ $=m_{t}+\xi\left(\frac{\partial\Pi_{P}}{\partial m}-\lambda\frac{\partial\text{Trust}}{\partial m}\right)$ (25)
where $\eta$ and $\xi$ are learning rates and $\lambda$ is the shadow price of user trust.
B.2 Model and Experimental Parameters
Table 3 provides a comprehensive list of the parameters used in the simulations.
Table 3: Model and Experimental Parameter Settings.
Category Parameter Name Symbol / Field Value Range or Default
General Max Steps max_steps 150 (Exp 1), 1000 (Exp 2)
Random Seed - 42
Agents Num Producers n_producers 80 (Exp 1), 100 (Exp 2)
HQ Productivity productivity_H Lognormal dist.
LQ Productivity productivity_L Lognormal dist.
Num Consumers n_consumers 200 (Exp 1), 300 (Exp 2)
Num Platforms n_platforms 1
Economic Platform Share $\theta$ / theta 0.25
Ad Revenue Base $\rho$ / rho 4.0
Labor Cost $w$ 8.0
AI Cost $r$ 1.0 (Baseline), swept in Exp 2
Technology HQ Elasticity $\sigma_{H}$ / sigma_H 0.75
LQ Elasticity $\sigma_{L}$ / sigma_L 1.5 (Baseline), swept in Exp 2
HQ Input Share $\delta_{H}$ / delta_H 0.35
LQ Input Share $\delta_{L}$ / delta_L 0.65
Behavioral Max Verification Cost $k_{max}$ 4.0
Risk Aversion risk_aversion Beta(2,3) dist.
Trust Decay Rate trust_decay 0.05
IPI Weights Pollution Weight w_pollution 0.35
Welfare Loss Weight w_welfare_loss 0.25
Trust Decay Weight w_trust_decay 0.25
Tech Risk Weight w_tech_risk 0.15
B.3 Overview of Experimental Procedures
The experimental validation was conducted through a series of structured simulation runs.
  • Baseline Experiment: The model was initialized and run for the maximum number of steps. Market states and agent behaviors were updated at each step, and key metrics were collected.
  • Exogenous Shock Experiment: Different types of shocks were injected at pre-set time steps (40, 70, 100, 130). The system’s response was measured by comparing the IPI trajectory before and after the shocks.
  • Weight Sensitivity Experiment: Multiple IPI weight combinations were constructed. The model was run for 100 steps for each combination, and the final correlation between IPI and social welfare was recorded.
  • Noise Robustness Experiment: Varying levels of measurement noise were added to the baseline model. Three independent trials were run for each noise level to calculate the volatility and measurement error of the IPI.
  • Cross-Platform Comparison: The network topology and agent preference parameters were altered to simulate different platform environments. Each configuration was run for 120 steps to compare final IPI and welfare outcomes.
  • Parameter Sweep & Policy Comparison (Exp 2): The simulation was run for 120 steps for each combination of AI cost ($r$) and LQ elasticity ($\sigma_{L}$). Additionally, six distinct policy configurations were simulated for 150 steps each to evaluate their impact on key outcomes.
B.4 Summary of Main Experimental Results
Tables 4 and 5 provide a consolidated summary of the key numerical findings from our two main experiments.
Table 4: Key Metrics from IPI Validation Experiment (Exp 1).
Phase Metric Value
Baseline Final IPI 0.611
Final Social Welfare (W) 67.78
IPI–W Correlation -0.839
Shock Resp. Avg. IPI Increase +37.5%
Recovery Rate (/step) 0.19–2.21
Weight Sens. Correlation Range 0.486–0.710
Best Weight Config. Equal
Noise Robust. Meas. Error (20% noise) 0.045
Avg. Volatility 0.033
Event Detect. Peak IPI Increase +21.6%
Recovery Rate -9.8%
Cross-Platform Min/Max IPI 0.527/0.630
Difference 0.103
Table 5: Key Results from Balanced Simulation (Exp 2).
Phase Metric Value
Baseline (1000 steps) Social Welfare 82.25
Pollution Density 0.607
IPI 0.624
Verification Rate 0.800
Trust Level 0.295
Parameter Sweep r–W Correlation +0.167
r–Pollution Correlation -0.770
Policy Comp. Best Policy (Joint) +2.3% W
Best Anti-Pollution Tech Interv.
B.7 Code and Data Availability
The complete Python scripts for Experiment 1 and 2 are available in the ‘main‘ branch of the following repository: github.com/Your-Repo/Information-Pollution-ABM
The required environment can be set up using Conda. The simulation results (CSV files and figures) are stored in the ‘results/‘ and ‘figures/‘ directories, respectively.
Appendix C Appendix C: Technical Details and Proofs
This appendix provides the technical derivations and formal proofs for the key results presented in the main text.
Table 6: Notation Summary
Symbol Description
$Q_{j}$ Content output of type $j\in\{H,L\}$ (high-/low-quality)
$A_{j}$ Total factor productivity for content type $j$
$K_{AI},L_{H}$ AI capital input and high-skilled labor input
$\delta_{j},\rho_{j}$ CES share and substitution parameters for type $j$
$\sigma_{j}=1/(1-\rho_{j})$ Elasticity of substitution for type $j$
$r,w$ Rental rate of AI capital and wage rate for labor
$c_{j}(r,w)$ Unit cost function for producing $Q_{j}$
$\theta$ Platform’s revenue share parameter
$\rho$ Ad‐revenue base per unit of amplified content
$m$ Moderation intensity chosen by platform
$\bm{\gamma}=(\gamma_{H},\gamma_{L})$ Algorithmic amplification weights
$k_{i}$ Verification cost of consumer $i$
$V$ Aggregate verification rate in the system
$\pi(s\!=\!H\mid q\!=\!H)$ Signal precision given pollution and verification
$W$ Social welfare function
$\mathrm{IPI}$ Information Pollution Index
C.1 Derivation of Asymmetric Production Costs
Derivation of the CES Cost Function.
For the CES production function $Q_{j}=A_{j}[\delta_{j}K_{AI}^{\rho_{j}}+(1-\delta_{j})L_{H}^{\rho_{j}}]^{1/\rho_{j}}$, the cost-minimization problem is:
$\min_{K_{AI},L_{H}}\quad$ $rK_{AI}+wL_{H}$ (26)
s.t. $A_{j}[\delta_{j}K_{AI}^{\rho_{j}}+(1-\delta_{j})L_{H}^{\rho_{j}}]^{1/\rho_{j}}=Q_{j}$
The Lagrangian is $\mathcal{L}=rK_{AI}+wL_{H}-\lambda(A_{j}[\dots]^{1/\rho_{j}}-Q_{j})$. The first-order conditions yield the optimal factor demand ratio:
$\frac{K_{AI}}{L_{H}}=\left(\frac{\delta_{j}}{1-\delta_{j}}\right)^{\sigma_{j}}\left(\frac{w}{r}\right)^{\sigma_{j}}$ (27)
Substituting this back into the production constraint and solving for the total cost yields the unit cost function:
$c_{j}(r,w)=\frac{1}{A_{j}}\left[\delta_{j}^{\sigma_{j}}r^{1-\sigma_{j}}+(1-\delta_{j})^{\sigma_{j}}w^{1-\sigma_{j}}\right]^{\frac{1}{1-\sigma_{j}}}$ (28)
Comparison of Logarithmic Elasticities.
The elasticity of the unit cost with respect to the AI capital price $r$ is equivalent to the cost share of AI capital, $s_{j,AI}$:
$\frac{\partial\log c_{j}}{\partial\log r}=\frac{\delta_{j}^{\sigma_{j}}r^{1-\sigma_{j}}}{\delta_{j}^{\sigma_{j}}r^{1-\sigma_{j}}+(1-\delta_{j})^{\sigma_{j}}w^{1-\sigma_{j}}}=s_{j,AI}$ (29)
To prove Proposition 1, we must show that $|s_{L,AI}|>|s_{H,AI}|$. Given Assumption 1 ($\sigma_{L}>1>\sigma_{H}$), it follows that $1-\sigma_{L}<0$ and $1-\sigma_{H}>0$. As $r$ decreases, the term $r^{1-\sigma_{L}}$ increases, while $r^{1-\sigma_{H}}$ decreases. This implies that the cost share of AI capital, $s_{L,AI}$, is more sensitive to changes in $r$ than $s_{H,AI}$. Therefore, the cost-reducing effect is larger for low-quality content. ∎
C.2 Proof Sketch for Existence of Equilibrium
The proof proceeds by backward induction.
  1. 1.
    Consumer Stage: For a given pollution density $\rho^{\prime}$, we define a mapping $T:[0,1]\to[0,1]$ where $T(V_{e})=F(k^{*}(\pi(\rho^{\prime},V_{e}),\rho^{\prime}))$ maps an expected verification rate $V_{e}$ to the actual rate. Since $k^{*}$ and $\pi$ are continuous in their arguments and $F(\cdot)$ is a continuous distribution function, $T$ is a continuous mapping from a compact, convex set to itself. By Brouwer’s Fixed-Point Theorem, a fixed point $V^{*}$ exists.
  2. 2.
    Producer Stage: Given producer heterogeneity, the aggregate supply function $Q_{j}^{S}(\gamma_{H},\gamma_{L})=\int q_{j,i}^{*}(\bm{\gamma})di$ is continuous in $\bm{\gamma}$ by the Theorem of the Maximum.
  3. 3.
    Platform Stage: The platform maximizes a continuous profit function $\Pi_{P}(m,\bm{\gamma})$ over a compact strategy space $\mathcal{S}=[0,1]\times[0,\bar{\gamma}]^{2}$. By the Weierstrass Extreme Value Theorem, a maximum exists.
The existence of optimal strategies in each stage implies the existence of an SPNE. ∎
C.3 Proof Sketches for Inefficiency and Comparative Statics
Proof of Theorem 2 (Threefold Market Failure).
The social planner’s problem is to choose $(Q_{H},Q_{L},m,V)$ to maximize social welfare. The first-order conditions (FOCs) are derived. We then compare these social FOCs with the FOCs from the decentralized equilibrium:
  • Production Externality: The producer’s FOC, $(1-\theta)\rho\gamma_{j}=MC_{j}$, lacks the social harm terms present in the planner’s FOC for $Q_{L}$.
  • Platform Failure: The platform’s FOCs for $m$ and $\gamma_{L}$ are based on maximizing private profit from engagement, which structurally deviates from maximizing social welfare.
  • Commons Externality: The consumer’s FOC for verification, $k_{i}=\text{Private Benefit}$, lacks the positive externality term $\partial\pi/\partial V$ that appears in the planner’s FOC for $V$.
These three wedges prove the Pareto inefficiency of the equilibrium. ∎
Proof of Proposition 1 (Paradox of AI Progress).
We use the Implicit Function Theorem on the system of equations defining the equilibrium.
  1. 1.
    The total derivative of the producer’s FOC with respect to $r$ shows a direct negative cost effect on $Q_{L}^{S}$.
  2. 2.
    The platform’s optimal response to this supply shift is to adjust $(\gamma_{L}^{*},m^{*})$, which further amplifies the initial shock. The combined effect leads to $\partial Q_{L}^{*}/\partial r<0$.
  3. 3.
    Since $Q_{L}^{*}$ increases more sensitively than $Q_{H}^{*}$ with a fall in $r$ (due to cost asymmetry), the pollution density $\rho^{\prime*}$ increases, so $\partial\rho^{\prime*}/\partial r<0$.
  4. 4.
    By the Envelope Theorem, $\partial W^{*}/\partial r=\partial\mathcal{L}/\partial r>0$, where $\mathcal{L}$ is the Lagrangian for the planner’s problem evaluated at the decentralized equilibrium. Since welfare $W^{*}$ is decreasing in pollution $\rho^{\prime*}$, and $\rho^{\prime*}$ is decreasing in $r$, welfare must be increasing in $r$. ∎